Suppose U(X,Y)=X^(2/3)*Y^(1/3), Px=3, Py=5. Find the Engel curves for goods X and Y and determine whether they are normal or inferior.
Answer #1

UE x=3, Py=5 Income = duidx durdy MRS= > M= Px x + Py. 4 M= 3xts M = 3x+ 5 | M = 9x engel for X - M= = Px x+ly.y 3x +sů M M=